a printing comapny sells small packages of personalized stationery for 7$ each, medium packages for 12$, and large packages for 15$ each. yesterday,the company sold 9 packages of stationary , collecting a total of $86. three times as many medium packages were sold large packages.

a.let s represent the number of small packages , m the number of medium packages, and l the number of large packages. write a system of three equations that represents the number of packages sold.
b.find the number of each size packages sold.

Answer by Nate(3500)

s + m + l = 9
7s + 12m + 15l = 86
m = 3l
Substitute:
s + 3l + l = 9
s + 4l = 9 which can later be substituted as s = 9 - 4l
Substitute:
7s + 12(3l) + 15l = 86
7s + 36l + 15l = 86
7s + 51l = 86
Substitute:
7(9 - 4l) + 51l = 86
63 - 28l + 51l = 86
23l = 23
l = 1
Remember:
m = 3l
m = 3
Remember:
s + m + l = 9
s + 3 + 1 = 9
s = 5